At 650°C phosphene vapor decomposes as follows:


APH3 → P4 + 6H₂ -rph (10h-¹)Cph,What size of plug flow reactor operating at 650°C and 11.4atm.is needed for conversion of 10 mole/hr of phosphine in a feed of 75% is phosphane and rest is inert


1
Expert's answer
2021-09-03T01:59:48-0400

Given FA0=10mole/hr, K=10/hr, XA=.75

Temp = 650oC=923oK,

P=11.4atm=11.4×101325=1155105pa

V=FA0/KCA0[(1+€A)ln1/(1-XA)-€AXA]----1

CA0=PA0/RT=

1155105pa/8.314pa.(m3/mol.K)(923K)

CAo= 150.52mol/m3

Put CAO in the above (1)

V=10/10×150.52[(1+0.75)ln(1/10.75)- 0.75×0.75]------(2)

On solving the above equation (2)

We get

V=0.01732 m3

Size of plug flow reactor is V = 0.01732m3